OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m= -7, (4,8) The equation of the line is y=

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

slope is -7 point is (4,8) There is a formula, y - y1 = m(x - x1) y - 8 = -7(x - 4) y - 8 = -7x + 28 y = -7x + 28 + 8 y = -7x + 36